Please help me figure out this problem.
Solve for Y
4y+5=4y-2
If there is more than one solution, seperate them with commas. if there is no solution, please put "no solution"

Answer by checkley79(3341)

4y+5=4y-2
4y-4y=-2-5
0y=-7
y=-7/0
y is undefined thus no solutions.